OpenStudy (anonymous):

All hats at a department store are on sale for 2/3 off. What would be the sale price on a hat that originally cost $12?

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

1-2/3=1/3 hence 1/3*12=4